### v1

@@ 5 Experiment
We compare BiKC with state-of-the-art (SOTA) imitation learning baselines. ACT [33] formulates the policy as a conditional variational autoencoder (cVAE), predicting a sequence of actions based on current observation. It adopts temporal ensemble to enhance action smoothness. However, ACT is designed to capture the primary behavior style, lacking the capability to handle multi-modality. DP [1] learns the policy as a diffusion model [11, 28], which denoises an action sequence iteratively from random noise, conditioned on historical observations. The action generation requires iterative network forward pass, resulting in substantial inference latency. We also extend DP to a keypose-conditioned variant called KDP for ablation study. Additionally, CP (Consistency Policy) is designed as a non-keypose version to serve as an ablation baseline. Here, the policy is formulated as a consistency model [27, 26] and learned from scratch by consistency training loss in Equation 3. CP predicts an action sequence conditioned on historical observations, which can be achieved through a single-step inference.
@@ 5.1 Bimanual Simulation
Tasks and Settings.
We evaluate BiKC on two bimanual tasks based on MuJoCo [30], Transfer and Insertion, developed by [33]. Both tasks involves multiple stages and find-grained coordination between two arms, as illustrated in Figure 4a. For Transfer task, four keyposes (i.e., start, grasp, transfer, end) are extracted as boundaries of three sub-stages, including: 1) right arm moves to fetch the cube, 2) the cube is grasped and lifted, and 3) the cube is handed over from the right to the left. Insertion also involves three stages separated by four keyposes (i.e., start, grasp, contact, insert), including: both arms 1) grasp the peg and socket, 2) lift peg and socket to allow contact, and 3) finally finish the bimanual insertion. The simulation environment includes two Interbotix vx300s robotic arm on both sides of a table-top. As shown in Figure 4b, the observation ${\bm{o}}_{t}$ consists of a $480\times 640$ RGB image rendered by MuJoCo from a static top-down view, as well as a 14D vector of current joint positions for both arms. The action is the expected joint positions of the two arms for the next timestep, which is also a 14D vector. The initial location of the cube, the socket and the peg are all randomly generated within a rectangle area in both demonstrations and evaluation rollouts. We refer to [33] for specific details. We use the 50 demonstrations open-sourced by [33] to train all models for each task. They are collected using a scripted policy at a control frequency of 50Hz, with each trajectory lasting 400 timesteps (i.e., 8 seconds). In the simulation experiments, we primarily focus on the impact of keypose on the success rate, while we disregard the evaluation of inference latency in the simulated tasks as the simulation is paused during policy inference.
@@
Task Sub-stage ACT     DP | KDP     CP | BiKC
Transfer To grasp 97 96 | 96 96 | 96
Lifting 90 96 | 96 93 | 95
Transferring 86 96 | 96 93 | 95
Insertion To grasp 93 83 | 95 83 | 98
To contact 90 47 | 89 53 | 96
To insert 32 37 | 48 38 | 43
Results.
We report the average success rate of each sub-stage in both simulated tasks in Table 1. All results are averaged across 3 seeds and 50 rollouts. Note that the sub-stages are defined same as [33] so we directly report results of ACT in the original paper. The top-two results are highlighted, which are predominantly achieved by the keypose-conditioned policies, namely KDP and BiKC. Particularly, for the Insertion task, BiKC significantly outperforms CP, with improvements of 16%, 43%, and 5% for each sub-task, respectively. Similarly, the ablation baseline KDP exhibits superior performance compared to DP, with success improvements of 12%, 42%, and 11% at each stage. Although ACT achieves highest success rates in some sub-tasks, its overall success rates for whole tasks are inferior. We suppose that it is because the third stage of both tasks involves fine-grained cooperation between two arms and are sensitive to operation accuracy. In other words, small errors will lead to collision and task failure. With the guidance from keyposes, KDP and BiKC are aware of the fine-grained sub-target and finally reaches higher success rate.
@@ 5.2 Real-World Experiments
We evaluate BiKC on three tasks involving multiple stages and bimanual coordination: Screwdriver Packing, Placing and Picking on Conveyor, and Cup Insertion. The Screwdriver Packing task aims to demonstrate BiKC’s ability to enhance sub-stage success rate while improving operational efficiency. The Placing and Picking on Conveyor task features a dynamic object, allowing us to emphasize the advantage of CM over DP in terms of inference speed. Finally, we assert that BiKC can capture multi-modality in Cup Insertion task, overcoming the limitations of ACT.
@@
Robot Platform.
We run real-world experiments with a low-cost open-source hardware system for bimanual teleoperation (ALOHA) [33]. ALOHA is comprised of two 7DOF master arms, two 7DOF puppet arms and accessories, as depicted in Figure 5. During teleoperation for recording demonstrations, a human manipulates the master arms, with their joint positions stored as action ${\bm{a}}_{t}$ and transmitted to the puppet arms for synchronization and execution. During deployment, the expected joint positions of puppet arms are provided by neural policies for autonomous execution. Four cameras installed in ALOHA offers visual observations. Two are positioned on the front and at the top of the frame, providing a static overall view of the workspace. The other two are mounted on the wrists of the puppets, offering closed views for fine-grained manipulation.
Tasks and Settings.
As shown in Figure 6, we conduct three multi-stage bimanual tasks with ALOHA, each of which focuses on a specific point. The first task, Screwdriver Packing, is a long-horizon task involving multiple challenging stages, such as eye-hand coordination (e.g., picking), high precision (e.g., insertion), bimanual coordination (e.g., flipping and pressing) as well as the transparent objects (e.g., the box cover). Due to the long horizon, this task is susceptible to the accumulation of error and latency along multiple stages, making it suitable for evaluating whether BiKC offers improved reliability and efficiency. The second task, Placing and Picking on Conveyor, involves moving objects across two stages. Initially, the left arm picks a bag and places it on the conveyor belt moving at 74 mm/s. As the bag approaches the right arm via the conveyor, the right arm picks and lifts the bag when it reaches an appropriate position. Since the bag is moving, if the right arm’s decision and execution are delayed by neural net inference, the bag may miss the suitable location, leading to task failure. We aim to use this task to highlight the advantages of BiKC in terms of inference speed. Additionally, we design a Cup Insertion task to insert a cup into a cup sleeve. We provide multi-modal demonstrations for this task: 1) The left arm first picks the sleeve, and then the right arm grasp the cup, and 2) The right arm first picks the cup, and then the left arm picks the sleeve, followed by insertion. We adopt this task to emphasize BiKC’s capabilities of multi-modal modeling. All these real-world tasks involve multiple stages and bimanual coordination, making them suitable for evaluating BiKC. Following the experimental setup outlined in previous work [33], we collected 50 demonstrations for each task.
@@
Results and Discussion.
We report the success rate, inference latency and total duration for three real-world experiments in Table 2. To quantify success rate for each sub-stage task, we follow the measurement approach adopted by [5]. For each stage, the success rate is calculated as Success divided by Attempts. Here, Attempts represents the number of successful cases from the previous stage, since the robot may fail and terminate at any stage. By employing this measurement, we can compute the final success rate for the entire task by taking the product of all sub-stage success rates. Additionally, the inference latency denotes the time required for the policy to output an action sequence. Operational efficiency is quantified by the total duration for finishing a fixed number of timesteps. All reported results represent an average of 20 rollouts, and top-2 results are highlighed in Table 2.
In the Screwdriver Packing task, BiKC outperforms CP by 35% in success rate, with the guidance from predicted keyposes. With additional condition on keyposes, BiKC is more aware of the sub-goal to be reached of short-horizon trajectories. This can be qualitatively verified by the differences in completion of the last stage of pressing. As shown in Figure 7b, CP tends to just move the grippers down slightly and fail to close the box, while BiKC is able to finish the stage substantially. While BiKC performs similarly as ACT in success rate and inference latency, its total operation duration (25.5s) is shorter than ACT by 11.5s, indicating superior operational efficiency. We suppose that the temporal ensembling technique, a key feature of ACT for high success rates and smooth motion, incurs additional time costs beyond policy inference. This experiment indicate BiKC’s superior performance in terms of success rate and operational efficiency, stemming from the keypose design and CM formulation. More discussions can be found in Appendix C.
In the Placing and Picking on Conveyor task, both ACT and DP variants fail to stably pick up the bag due to excessive inference latency, showing lack of ability to track dynamic objects. In contrast, CP and BiKC achieve a success rate of 100%, attributed to their policy formulation as consistency models, which generate action sequences by one-step inference with few latency. This experiment, involving dynamic factors, underscores the advantages of CM and the importance of inference efficiency as indicated in Figure 7a.
@@ 6 Conclusion
Limitations and Future Work.
We believe that this work can be extended from both learning and representation perspectives. Regarding the learning aspect, despite of accelerated trajectory generation, samples generated by CM are not as smooth as those by ACT. A similar phenomenon also occurs in image synthesis in the original work [27, 26], where the details in images generated by CM are compromised without high fidelity. Although this can be mitigated by the temporal ensemble proposed in [33], the rationality of performing IL on short-horizon trajectories should be further verified due to the existence of modern motion planning and control approaches. It is promising to explore whether it will be efficient and successful that IL is applied only in a coarser manner such as high-level plans and geometric constraints extracted from demonstrations. By doing so, it becomes possible to guarantee reaching keyposes or satisfying constraints, thereby further ensuring that the completion of sub-stages. Regarding the representation aspect, we utilize joint positions as keyposes and sub-goals in this work because it aligns with the recommended state and action form of real robots, ensuring coherence in our representation. However, it is important to note that joint representations may not be as intuitive and generalized as desired, as they require forward kinematics to transform joint configurations into meaningful SE(3) poses. Furthermore, despite their popularity, we discovered that both joint positions and SE(3) poses are inherently robot-centric and may not accurately delineate the boundaries of sub-stages in practical scenarios, as they primarily reflect the states of the robots themselves rather than capturing the task context. The completion of manipulation tasks usually depends on desired state transition of objects instead of robots [6]. Hence, it is necessary to leverage a representation that captures changes in object-centric states and further the whole workspace, and take them as sub-goals of motion. This will be more helpful for autonomous agents to understand task progress and achieve sub-goals. A scene graph is a promising formulation because it reveals the high-level relationships among objects in the workspace and can be extracted by advanced vision models [13, 14, 35].
@@
[…]
The authors have no competing interests to declare that are relevant to the content of this article.
@@ Appendix A Recapping Designs of Consistency Models
Discretization curriculum
The continuous noise level $[{\epsilon},\sigma_{\max}]$ is discretized into $N(k)$ points at the $k$-th training iteration, with $K$ iterations in total and $N(k)$ growing as $k$ increases. The new curriculum adopts a step-shaped curriculum where $N(k)$ doubles after a fixed number of iterations. Specifically,
@@
Noise schedule
Although [26] propose to sample $\sigma_{i}$ according to a lognormal distribution from the discretized sequence, we find that the original uniform distribution sampling performs better in our tasks. Therefore we stick to the old choice where $i\sim{\mathcal{U}}\{1,2,\cdots,N(k)-1\}$.
Parameterization
In according to satisfy the self-consistency condition mentioned in Section 3, authors of CM [27] leverage a special parameterization to make ${\bm{f}}_{\bm{\theta}}({\mathbf{x}}_{\epsilon},{\epsilon})={\mathbf{x}}_{\epsilon}$ hold by design:

### v2

@@ 5 Experiment
We compare BiKC with state-of-the-art (SOTA) imitation learning baselines. ACT [37] formulates the policy as a conditional variational autoencoder (cVAE), predicting a sequence of actions based on current observation. It adopts temporal ensemble to enhance action smoothness. However, ACT is designed to capture the primary behavior style, lacking the capability to handle multi-modality. DP [2] learns the policy as a diffusion model [13, 31], which denoises an action sequence iteratively from random noise, conditioned on historical observations. The action generation requires iterative network forward pass, resulting in substantial inference latency. Additionally, CP (Consistency Policy) is designed as a non-keypose version to serve as an ablation baseline. Here, the policy is formulated as a consistency model [30, 29] and learned from scratch by consistency training loss in Equation 3. CP predicts an action sequence conditioned on historical observations, which can be achieved through a single-step inference.
@@ 5.1 Bimanual Simulation
Tasks and Settings. We evaluate BiKC on two bimanual tasks based on MuJoCo [33], Transfer and Insertion, developed by [37]. Both tasks involves multiple stages and find-grained coordination between two arms, as illustrated in Figure 4a. For Transfer task, four keyposes (i.e., start, grasp, transfer, end) are extracted as boundaries of three sub-stages, including: 1) right arm moves to fetch the cube, 2) the cube is grasped and lifted, and 3) the cube is handed over from the right to the left. Insertion also involves three stages separated by four keyposes (i.e., start, grasp, contact, insert), including: both arms 1) grasp the peg and socket, 2) lift peg and socket to allow contact, and 3) finally finish the bimanual insertion. The simulation environment includes two Interbotix vx300s robotic arm on both sides of a table-top. As shown in Figure 4b, the observation ${\bm{o}}_{t}$ consists of a $480\times 640$ RGB image rendered by MuJoCo from a static top-down view, as well as a 14D vector of current joint positions for both arms. The action is the expected joint positions of the two arms for the next timestep, which is also a 14D vector. The initial location of the cube, the socket and the peg are all randomly generated within a rectangle area in both demonstrations and evaluation rollouts. We refer to [37] for specific details. We use the 50 demonstrations open-sourced by [37] to train all models for each task. They are collected using a scripted policy at a control frequency of 50Hz, with each trajectory lasting 400 timesteps (i.e., 8 seconds). In the simulation experiments, we primarily focus on the impact of keypose on the success rate, while we disregard the evaluation of inference latency in the simulated tasks as the simulation is paused during policy inference.
@@
Task Sub-stage ACT DP     CP | BiKC
Transfer To grasp 97 96 96 | 96
Lifting 90 96 93 | 95
Transferring 86 96 93 | 95
Insertion To grasp 93 83 83 | 98
To contact 90 47 53 | 96
To insert 32 37 38 | 43
Results. We report the average success rate of each sub-stage in both simulated tasks in Table 1. All results are averaged across 3 seeds and 50 rollouts. Note that the sub-stages are defined same as [37] so we directly report results of ACT in the original paper. The top-two results are highlighted, which are predominantly achieved by BiKC. Particularly, for the Insertion task, BiKC significantly outperforms CP, with improvements of 16%, 43%, and 5% for each sub-task, respectively. Although ACT achieves highest success rates in some sub-tasks, its overall success rates for whole tasks are inferior. We suppose that it is because the third stage of both tasks involves fine-grained cooperation between two arms and is sensitive to operation accuracy. In other words, small errors will lead to collision and task failure. With guidance from keyposes, BiKC is aware of the fine-grained sub-target and finally reaches higher success rates.
@@ 5.2 Real-World Experiments
We evaluate BiKC on three real-world tasks featuring multiple sub-stages, intricate coordination, and rich contact: Screwdriver Packing, Placing and Picking on Conveyor, and Cup Insertion. The Screwdriver Packing task aims to demonstrate BiKC’s ability to enhance sub-stage success rate while improving operational efficiency. The Placing and Picking on Conveyor task features a dynamic object, allowing us to emphasize the advantage of CM over DP in terms of inference speed. Finally, we assert that BiKC can capture multi-modality in Cup Insertion task, overcoming the limitations of ACT.
@@
Robot Platform. We run real-world experiments with a low-cost open-source hardware system for bimanual teleoperation (ALOHA) [37]. ALOHA is comprised of two 7DOF master arms, two 7DOF puppet arms and accessories, as depicted in Figure 5. During teleoperation for recording demonstrations, a human manipulates the master arms, with their joint positions stored as action ${\bm{a}}_{t}$ and transmitted to the puppet arms for synchronization and execution. During deployment, the expected joint positions of puppet arms are provided by neural policies for autonomous execution. Four cameras installed in ALOHA offers visual observations. Two are positioned on the front and at the top of the frame, providing a static overall view of the workspace. The other two are mounted on the wrists of the puppets, offering closed views for fine-grained manipulation.
Tasks and Settings. As shown in Figure 6, we conduct three multi-stage bimanual tasks with ALOHA, each of which focuses on a specific point. The first task, Screwdriver Packing, is a long-horizon task involving multiple challenging stages, such as eye-hand coordination (e.g., picking), high precision (e.g., insertion), bimanual coordination (e.g., flipping and pressing) as well as the transparent objects (e.g., the box cover). Due to the long horizon, this task is susceptible to the accumulation of error and latency along multiple stages, making it suitable for evaluating whether BiKC offers improved reliability and efficiency. The second task, Placing and Picking on Conveyor, involves moving objects across two stages. Initially, the left arm picks a bag and places it on the conveyor belt moving at 74 mm/s. As the bag approaches the right arm via the conveyor, the right arm picks and lifts the bag when it reaches an appropriate position. Since the bag is moving, if the right arm’s decision and execution are delayed by neural net inference, the bag may miss the suitable location, leading to task failure. We aim to use this task to highlight the advantages of BiKC in terms of inference speed. Additionally, we design a Cup Insertion task to insert a cup into a cup sleeve. We provide multi-modal demonstrations for this task: 1) The left arm first picks the sleeve, and then the right arm grasp the cup, and 2) The right arm first picks the cup, and then the left arm picks the sleeve, followed by insertion. We adopt this task to emphasize BiKC’s capabilities of multi-modal modeling. All these real-world tasks involve multiple stages and bimanual coordination, making them suitable for evaluating BiKC. Following the experimental setup outlined in previous work [37], we collected 50 demonstrations for each task. Throughout all demonstrations and rollouts, the initial workspace layout remains constant.
@@
Results and Discussion. We report the success rate, inference latency and total duration for three real-world experiments in Table 2. To quantify success rate for each sub-stage task, we follow the measurement approach adopted by [6]. For each stage, the success rate is calculated as Success divided by Attempts. Here, Attempts represents the number of successful cases from the previous stage, since the robot may fail and terminate at any stage. By employing this measurement, we can compute the final success rate for the entire task by taking the product of all sub-stage success rates. Additionally, the inference latency denotes the time required for the policy to output an action sequence. Operational efficiency is quantified by the total duration for finishing a fixed number of timesteps. All reported results represent an average of 20 rollouts, which is a widely adopted protocol in this field of research [37, 2]. Top-2 results are highlighed in Table 2.
In the Screwdriver Packing task, BiKC outperforms CP by 35% in success rate, with the guidance from predicted keyposes. With additional condition on keyposes, BiKC is more aware of the sub-goal to be reached of short-horizon trajectories. This can be qualitatively verified by the differences in completion of the last stage of pressing. As shown in Figure 7b, CP tends to just move the grippers down slightly and fail to close the box, while BiKC is able to finish the stage substantially. While BiKC performs similarly as ACT in success rate and inference latency, its total operation duration (25.5s) is shorter than ACT by 11.5s, indicating superior operational efficiency. We suppose that the temporal ensembling (TE) technique, a key feature of ACT for high success rates and smooth motion, incurs additional time costs beyond policy inference. This experiment indicate BiKC’s superior performance in terms of success rate and operational efficiency, stemming from the keypose design and CM formulation. More discussions can be found in Appendix C.
In the Placing and Picking on Conveyor task, both ACT and DP variants fail to stably pick up the bag due to excessive inference latency, showing lack of ability to track dynamic objects. In contrast, CP and BiKC achieve a success rate of 100%, attributed to their policy formulation as consistency models, which generate action sequences by one-step inference with few latency. This experiment, involving dynamic factors, underscores the advantages of CM and the importance of inference efficiency as indicated in Figure 7a. Additionally, we observe that the inference latency of ACT is comparable to that of CP and BiKC. However, ACT fails to catch moving objects in a timely manner, which is attributed to TE again. Specifically, ACT predicts an action sequence at each timestep and TE performs a weighted average over historical predictions, significantly increasing decision delay. In contrast, CP and BiKC only need to perform inference every $H$ steps, enabling latency amortization and real-time reactions.
@@ 6 Conclusion
Limitations and Future Work. We believe that the framework can be extended from learning and representation perspectives. Regarding learning, despite of accelerated trajectory generation, samples generated by CM are not as smooth as those by ACT. A similar phenomenon also occurs in image synthesis in the original work [30, 29], where visual details generated by CM are compromised without high fidelity. Although this can be mitigated by the TE trick [37], the necessity of IL on short-horizon trajectories should be further verified due to the existence of modern motion planning and control approaches. It is promising to explore whether it will be efficient and successful that IL is applied only in a coarser manner such as high-level plans and geometric constraints extracted from demonstrations. By doing so, it becomes possible to guarantee reaching keyposes or satisfying constraints, thereby further ensuring completing sub-stages. Regarding representation, we utilize joint positions as keyposes and sub-goals in this work because it aligns with the recommended state and action form of real robots, ensuring coherence in our representation. However, it is important to note that joint representations may not be as intuitive and generalized as desired, as they require forward kinematics to transform joint configurations into meaningful SE(3) poses. Furthermore, despite their popularity, we discovered that both joint positions and SE(3) poses are inherently robot-centric and may not accurately delineate the boundaries of sub-stages, as they primarily reflect the states of the robots themselves rather than capturing the task context. The completion of manipulation tasks usually depends on desired state transition of objects instead of robots [7]. Hence, it is necessary to leverage a representation that captures changes in object-centric states and further the whole workspace, and take them as sub-goals of motion. This will be more helpful for autonomous agents to understand task progress and achieve sub-goals. A scene graph is a promising formulation because it reveals the high-level relationships among objects in the workspace and can be extracted by advanced vision models [15, 16, 39].
The keypose identification approach warrants further investigation. While the proposed heuristics address a range of tasks, certain situations remain out of scope. A significant challenge arises when human demonstrations are suboptimal or contain retries, a common issue in IL. In such cases, momentary pauses by human operator may be erroneously identified as keyposes, despite their irrelevance to task completion or sub-stage transitions.nother challenge emerges when sub-task boundaries lack clear definition, as exemplified by periodic tasks such as rotating a faucet handle. In such scenarios, delineating the task using a limited number of keyposes is challenging, as cyclic trajectories cannot be accurately interpolated from a finite set of intermediate steps. Future work should investigate these challenged cases in greater depth.
@@
6.0.1 Acknowledgements
This study is partially supported by the Innovation and Technology Commission of the HKSAR Government under the InnoHK initiative as well as Research Grant Council (Ref: 17200924).
@@ Appendix A Recapping Designs of Consistency Models
Discretization curriculum The continuous noise level $[{\epsilon},\sigma_{\max}]$ is discretized into $N(k)$ points at the $k$-th training iteration, with $K$ iterations in total and $N(k)$ growing as $k$ increases. The new curriculum adopts a step-shaped curriculum where $N(k)$ doubles after a fixed number of iterations. Specifically,
@@
Noise schedule Although [29] propose to sample $\sigma_{i}$ according to a lognormal distribution from the discretized sequence, we find that the original uniform distribution sampling performs better in our tasks. Therefore we stick to the old choice where $i\sim{\mathcal{U}}\{1,2,\cdots,N(k)-1\}$.
Parameterization In according to satisfy the self-consistency condition mentioned in Section 3, authors of CM [30] leverage a special parameterization to make ${\bm{f}}_{\bm{\theta}}({\mathbf{x}}_{\epsilon},{\epsilon})={\mathbf{x}}_{\epsilon}$ hold by design: